OpenStudy (anonymous):

HELP!! Object A with a mass of 1.5 kilograms is moving with a velocity of +11.2 meters/second (moving in the +x direction). It has a perfectly elastic collision with stationary object B that has a mass of 4.5 kilograms. After the collision, object B travels with a velocity of +5.6 meters/second. What is the final velocity for object A? I'll give out METALS to the right answer!!!

OpenStudy (anonymous):

This is a "momentum" question. That is, energy is conserved after impact. The idea here is that the total momentum in the objects is the same before and after. The equation you set up is: M1V1 + M2V2 = M1(V1)' + M2(V2)' Givens: M1 = 1.5kg M2 = 4.5kg V1 = 11.2 m/s V2 = 0 m/s (V2)' (after impact) = 5.6 m/s Subbing into the equation above, we get: (1.5)(11.2) + (4.5)(0) = (1.5)(x) + (4.5)(5.6) Rearrange and calculate: x = -5.6 What this means is, object 1 is moving 5.6 m/s in the opposite direction now. Since we have 100% elastic conditions, there is no energy absorbed by the objects. There is your answer.